======20241130====== =====Time===== 1630 - 2300 =====Attendance===== Connor, Ian, Tara =====Spaces Occupied===== 401, 433C, Dome =====Goals===== =====Equipment Used===== ^ Scope ^ Camera ^ Focus ^ |Main|PICSI2 w/ 178|2618.7| |Finder|ASI 482|2465| =====Frame Table===== ^ Time Observed ^ Source Name ^ Filter ^ Int Time [s] ^ Gain ^ Airmass ^ Mag (filter) ^ RA [h] ^ DEC [deg] ^ Source Type ^Scope ^ Purpose ^ Notes^ | 20241130T1715 | Venus | UGI | 100ms | 150 | airmass | -4.05(V) | 19 | 39 | Planet | Main |Pol target | tracking? Others?| | 20241130T1757 | 4C56.27 | UGI | | 150 | airmass | 15(V) | 18 | 56 | Blazar | Main |Pol target | tracking? Others?| | 20241130T1822 | 4C27.50 | UGI | | 150 | airmass | 15(V) | 23 | 27 | Quazar | Main |Pol target | tracking? Others?| | 20241130T1838 | S4 1749+70 | UGI | | 150 | airmass | 16(V) | 17 | 70 | Blazar | Main |Pol target | tracking? Others?| | 20241130T1900 | PKS B2258-022| UGI | | 150 | airmass | 18(V) | 23| -01 | Quazar | Main |Pol target | tracking? Others?| | 20241130T1928 | BL Lac| R | 15s| 150 | airmass | 14(V) | 22| 42 | Blazar | Main |Pol target | tracking? Others?| | 20241130T2024 | HD 19820| RG | 1s| 150 | airmass | 7.1(V) | 03| 59 | Pol star | Main |Pol verification | tracking? Others?| | 20241130T2032 | HD 204827| RG | 1s| 150 | airmass | 8.35(V) | 21| 58 | Pol star | Main |Pol verification | tracking? Others?| | 20241130T2047 | HD212311| RG | 1s| 150 | airmass | 8.12(V) | 22| 56 | UnPol star | Main |Pol verification | tracking? Others?| =====Log===== - Regular startup procedures - North mirror door not opening with command; East and West also struggled because of this - Tried closing and reopening them twice, north worked on second try - BUT East decided to stop opening halfway - Closed and reopened and then heard a bang followed by harmonic clanging - East mirror door had fallen down back over the mirror - Clevis pin has disappeared - Checked mirror; No visible damage and pin does not appear to be there - Replaced Clevis pin - SUCCESS; All doors open - Finished startup procedures - Followed Ian/Dr. Meyer's observing plan with their oversight - Attempted flats with new string lights - No glare but indeed too dim - Used regular in-dome lights and the yellow LED lamp to boost flux - Closing procedures in dome to prep for darks - North mirror door didn't budge - Helped it along by hand, success - Took darks - Closed up the control room =====ToDo===== =====Notes===== =====Weather===== =====Data===== {{tag>[picsarr]}}